1) Find the equations of the staright line through the points (1,2,3) and parallel to the straight line joining the points (-4,7,2) and (5,-3,-2) . 2) Show that the equation of the projection of the staright line (x-1)/2 = (y+1)/-1 = (z-3) /4 on the plane x+2y + z = 6 are (x-3)/4. = (y+2)/-7 = (z-7)/10

1)  Find the equations of the staright line through the points (1,2,3) and parallel to the straight line joining the points (-4,7,2) and (5,-3,-2) . 
  
2)  Show that the equation of the projection of the staright line (x-1)/2 = (y+1)/-1 =  (z-3) /4 on the plane x+2y + z = 6 are (x-3)/4. = (y+2)/-7 = (z-7)/10

Grade:12

1 Answers

Susmita
5 years ago
Direction vector of line joining (-4,7,2) and (5,-3,-2)
is vec(d)=-9i+10j+4k.
Equation of line parallel to vec(d) and passing through (1,2,3) is
\frac{x-1}{-9}=\frac{y-2}{10}=\frac{z-3}{4}